An investor invested a total of $2,300 in two mutual funds. one fund earned a 9% profit while the other earned a 2% profit was $95. how much was invested in each mutual fund
Answer by mananth(16946)

Part I 9.00% per annum ------------- Amount invested =x
Part II 2.00% per annum ------------ Amount invested = y =
2300
Interest----- 95

Part I 9.00% per annum ---x
Part II 2.00% per annum ---y
Total investment
x + y= 2300 -------------1
Interest on both investments
9.00% x + 2.00% y= 95
Multiply by 100
9 x + 2 y= 9500.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -20700.00
Add this to (2)
0 x -7 y= -11200
divide by -7
y = 1600
Part I 9.00% $ 700
Part II 2.00% $ 1600

CHECK
700 --------- 9.00% ------- 63.00
1600 ------------- 2.00% ------- 32.00
Total -------------------- 95.00

m.ananth@hotmail.ca